find monthly payment on a 5year loan of 4500 at an annual interest rate of 6.5%
Answer by Phillips54(20)

Use cross multiplication...
5/4500=6.5/x
4500*6.5=29250
29250/5=5850
after 5 years with an interest rate of 6.5, will be 5850.